the energy value of carbohydrates is 4.0 cal/g. calculate the energy in 250 g of carbohydrates, and express the value in joules
Answer #1

I hope this helps you.

Energy in 1 gram of carbohydrates = 4 calories

Hence energy in 250 g of carbohydrates = 4 * 250 = 1000 calories

1 calorie = 4.184 joules

Hence, 1000 calories = 4.184 * 1000 = 4184 Joules

Do comment if you have any doubt.

Feedback appreciated..

All the best! :)